OpenStudy (anonymous):

Betsy wants to paint the walls of her bedroom. Her bedroom measures 8 ft. by 10 ft. and has 8 ft. ceilings. Not counting doors and windows, how many square feet of wall will Betsy need to paint? 288<---my answer 80 128 160 The perimeter of a rectangle is four more than the area of the rectangle. Find the area of the rectangle, if its width is 1 cm. 0.5 cm^2 1 cm^2 8 cm^2 2 cm^2

Directrix (directrix):

I'll crank out #1 on paper and see what I get.

OpenStudy (anonymous):

ok

Directrix (directrix):

Do you think that the ceiling counts as a wall? I guess not. Did you include the ceiling area.

OpenStudy (anonymous):

i did this 2(8*10+8*8) and i got 288

Directrix (directrix):

I did (8 + 8 + 10 + 10) = 36 to get the perimeter of the base and then multiplied by 8 to get 288

Directrix (directrix):

@kittymeow101 I think we're good on this problem.

Directrix (directrix):

Now, to #2.

OpenStudy (anonymous):

ok wb the next one?

Directrix (directrix):

Did you get an answer on #2?

Directrix (directrix):

Perimeter is (x + x + 1 + 1) = ? Area = x * 1 = x